A bike leaves @ 8:00 am traveling at 10 mph. A car leaves, heading in the same direction @ 11:00 am traveling at 50 mph. How many minutes after 11:00 will the car overpass the bike. The computer program gives the answer of 45 minutes, but I can't figure out how? Thanks!
Found 2 solutions by Nate, saltzy57:
Answer by Nate(3500)

'y' expresses distance traveled
y = 10mph + 30 the biker has already traveled 30 miles when the car leaves
y = 50mph
Combine:
10x + 30 = 50x
30 = 40x
3/4 = x
Sixty minutes in an hour. (60)(3/4) = 45 minutes

Answer by saltzy57(8)

ok after 3 hrs, the bike will have travelled 30 miles and now the car starts at 50 mph...we can write this equation as:
30 + 10x = 50x solving for x
x= 3/4 of an hour or 45 mins